# Efficient Error and Variance Estimation

### v3

@@ Abstract
Randomized matrix algorithms have become workhorse tools in scientific computing and machine learning. To use these algorithms safely in applications, they should be coupled with posterior error estimates to assess the quality of the output. To meet this need, this paper proposes two diagnostics: a leave-one-out error estimator for randomized low-rank approximations and a jackknife resampling method to estimate the variance of the output of a randomized matrix computation. Both of these diagnostics are rapid to compute for randomized low-rank approximation algorithms such as the randomized SVD and Nyström, and they provide useful information that can be used to assess the quality of the computed output and guide algorithmic parameter choices.
@@ 1 Introduction
In recent years, randomness has become an essential tool in the design of matrix algorithms [5, 11, 16, 31], with randomized algorithms proving especially effective for matrix low-rank approximation. Since the outputs of randomized matrix algorithms are stochastic quantities, these algorithms should be supported in practice with posterior error estimates and other quality metrics that inform the user about the accuracy of the computational output.
@@
By using novel downdating formulas (see Eqs. 15 and 17 below), we can rapidly compute both of these estimators for widely used low-rank approximation methods such as the randomized SVD and the randomized Nyström approximation. These methods are also sample-efficient, requiring no additional information beyond that used in the original algorithm. The speed and efficiency of these diagnostics make them compelling additions to workflows involving randomized matrix computation.
@@ 1.1 Leave-one-out error estimation
We begin by motivating our first diagnostic, a leave-one-out estimator for the error of a low-rank matrix approximation.
Suppose we want to approximate a positive-semidefinite (psd) matrix $\boldsymbol{A}\in\real^{d\times d}$, which we can only access by the matrix–vector product operation $\boldsymbol{\omega}\mapsto\boldsymbol{A}\boldsymbol{\omega}$. Using the matrix–vector product operation alone, we can compute the matrix–matrix product $\boldsymbol{A}\boldsymbol{\Phi}$ with a (random) matrix $\boldsymbol{\Phi}\in\real^{d\times s}$ with $s$ columns and form the Nyström approximation [9, 27, 30]
@@
where $\boldsymbol{\Omega}$ is populated with statistically independent standard Gaussian entries. The quality of the Nyström approximation improves with higher approximation rank $s$ and subspace iteration steps $q$. As we detail in Algorithm 1, in $\mathcal{O}(qsd^{2})$, we can compute $\boldsymbol{A}\langle\boldsymbol{\Phi}\rangle$ in the form of a compact eigenvalue decomposition:
@@
where $\boldsymbol{V}\in\real^{d\times s}$ has orthonormal columns and $\boldsymbol{\Lambda}\in\real_{+}^{s\times s}$ is diagonal.
@@
How do we estimate the error $\left\|\boldsymbol{A}-\boldsymbol{A}\langle\boldsymbol{\Phi}\rangle\right\|_{\rm F}$ of the Nyström approximation when we can only access $\boldsymbol{A}$ through matrix–vector products?
The leave-one-out error estimator provides exactly such an estimate. Moreover, this estimate is sample-efficient in that it requires no additional matrix–vector products with $\boldsymbol{A}$ beyond those used to form $\boldsymbol{A}\langle\boldsymbol{\Phi}\rangle$.
@@
$\boldsymbol{\Omega}\mapsto\boldsymbol{X}=\boldsymbol{X}(\boldsymbol{\Omega})=\boldsymbol{A}\langle\boldsymbol{A}^{q}\boldsymbol{\Omega}\rangle.$
Introduce replicates $\boldsymbol{X}^{(1)},\ldots,\boldsymbol{X}^{(s)}$ obtained by recomputing $\boldsymbol{X}$ with each column of $\boldsymbol{\Omega}$ left out in turn:
$\boldsymbol{X}^{(j)}=\boldsymbol{X}\big(\boldsymbol{\Omega}^{(j)}\big)\quad\text{for $j=1,2,\ldots,s$},$
where $\boldsymbol{\Omega}^{(j)}$ is $\boldsymbol{\Omega}$ without its $j$th column. Define the leave-one-out error estimator
$\widehat{\operatorname{Err}}^{2}(\boldsymbol{X},\boldsymbol{A})\coloneqq\frac{1}{s}\sum_{i=1}^{s}\big\|\big(\boldsymbol{A}-\boldsymbol{X}^{(j)}\big)\boldsymbol{\omega}_{j}\big\|^{2},$
where $\boldsymbol{\omega}_{j}$ denotes the $j$th column of $\boldsymbol{\Omega}$ and $\left\|\cdot\right\|$ is the Euclidean norm. As we show in more generality in Theorem 2.1, this error estimator is an unbiased estimator for the mean-square error of the rank-(s1)(s-1) Nyström approximation:
@@
This estimate is rapid to compute, requiring at most $\mathcal{O}(s^{2}d)$ operations (and only $\mathcal{O}(s^{3})$ operations if $q=0$). The error estimate $\widehat{\operatorname{Err}}(\boldsymbol{X},\boldsymbol{A})$ requires no information about $\boldsymbol{A}$ beyond what is required to form the approximation $\boldsymbol{X}$.
The leave-one-out error estimator is demonstrated in Fig. 1, which plots the mean error
@@
and leave-one-out error estimate $\widehat{\operatorname{Err}}(\boldsymbol{X},\boldsymbol{A})$ for a randomized Nyström approximation Eq. 1 of a psd kernel matrix $\boldsymbol{A}\in\real^{10^{4}\times 10^{4}}$ formed from a random subsample of $10^{4}$ points from the QM9 dataset [21, 22]. We choose $q=0$ and use $1000$ trials to estimate the average error $\operatorname{Err}(\boldsymbol{A},\boldsymbol{X})$ and to compute the mean and standard deviation of the error estimate $\widehat{\operatorname{Err}}(\boldsymbol{A},\boldsymbol{X})$. We see that the estimate $\widehat{\operatorname{Err}}(\boldsymbol{X},\boldsymbol{A})$ closely tracks the true error. Moreover, the error estimate $\widehat{\operatorname{Err}}(\boldsymbol{X},\boldsymbol{A})$ is fast to compute, with the error estimate taking less than 1% of the total runtime to form.
Figure 1: Leave-one-out error estimator. Average relative error $\operatorname{Err}(\boldsymbol{X},\boldsymbol{A})/\left\|\boldsymbol{A}\right\|_{\rm F}$ and normalized error estimate $\widehat{\operatorname{Err}}(\boldsymbol{X},\boldsymbol{A})/\left\|\boldsymbol{A}\right\|_{\rm F}$ for Nyström approximation $\boldsymbol{X}$ to a psd kernel matrix $\boldsymbol{A}\in\real^{10^{4}\times 10^{4}}$. Error and error estimate were computed by an average of $1000$ trials. Error bars show one standard deviation.
@@ 1.2.3 Jackknife variance estimation
Guarantees for this estimator are provided in Theorem 3.1. Algorithm 3 and LABEL:list:spectral_clustering provide pseudocode and a MATLAB implementation of Nyström-accelerated spectral clustering with the jackknife estimate $\operatorname{Jack}(\boldsymbol{X})$. The cost of forming the jackknife estimate $\operatorname{Jack}(\boldsymbol{X})$ is $\mathcal{O}(s^{2}d)$, much faster than the $\mathcal{O}(qsd^{2})$ cost of Nyström-accelerated spectral clustering.
@@ 1.2.4 Numerical example
the mean and standard deviation for the standard deviation estimate $\operatorname{Jack}(\boldsymbol{X})$, and the empirical success probability for spectral clustering.
Figure 2 shows the results. The estimate $\operatorname{Jack}(\boldsymbol{X})$ overestimates the standard deviation, but by a modest amount (a factor of five at most). The jackknife is not quantitatively sharp, but it is a reliable indicator of whether the variance is high or low.
The jackknife estimate $\operatorname{Jack}(\boldsymbol{X})$ can be used to determine whether the Nyström-based approximations to the eigenvectors are accurate enough for clustering task. At rank $s=50$, clustering is never performed correctly, and the jackknife estimate $\operatorname{Jack}(\boldsymbol{X})\approx 2$ is high. As the approximation rank $s$ is increased, clustering begins to succeed with higher and higher probability and the jackknife estimate $\operatorname{Jack}(\boldsymbol{X})$ decreases, indicating reduced variability. This observation suggests two possible uses for the jackknife estimate:
@@
Both of these uses demonstrate the potential for jackknife variance estimation to be helpful in incorporating randomized low-rank approximation into general-purpose software.
@@ 2.2 Alternatives
A second alternative is provided by randomized norm estimates [16, §§4–5]. For isotropic random vectors $\boldsymbol{\nu}_{1},\ldots,\boldsymbol{\nu}_{t}$ independent of the random approximation $\boldsymbol{X}$,
(9) $\frac{1}{t}\sum_{j=1}^{t}\left\|(\boldsymbol{A}-\boldsymbol{X})\boldsymbol{\nu}_{j}\right\|^{2}\quad\text{is an unbiased estimate for}\quad\left\|\boldsymbol{A}-\boldsymbol{X}\right\|_{\rm F}^{2}.$
The leave-one-out error estimator Eq. 8 has a similar form except that we reuse the test vectors $\boldsymbol{\omega}_{1},\ldots,\boldsymbol{\omega}_{s}$ instead of generating fresh vectors $\boldsymbol{\nu}_{1},\ldots,\boldsymbol{\nu}_{t}$. Because of this reuse, the leave-one-our error estimator is more sample-efficient than the norm estimate Eq. 9. For the leave-one-out error estimator, every matrix–vector product $\boldsymbol{A}\boldsymbol{\omega}_{j}$ is used to both form the approximation $\boldsymbol{X}$ and estimate the error, no extra matrix–vector products required.
@@ 3.5 Related work: bootstrap for randomized matrix algorithms
Bootstrap resampling [6, §5], a close relative of the jackknife, has seen several applications to matrix computations. An early use case was to provide confidence intervals for eigenvalues and eigenvectors of sample covariance matrices [25, §7.2].
A more recent line of work, led by Lopes and collaborators, applies bootstrap resampling to randomized matrix algorithms [13, 14, 15, 32]. The work closest to ours [13] uses the bootstrap to provide asymptotically sharp estimates of the error quantiles with regards to general error metrics for each singular value and singular vector computed by a linear sketched SVD algorithm. In this special case, the bootstrap provides more fine-grained and quantitatively sharp information than the matrix jackknife. Unfortunately, the linear SVD is a poor computational method because its error decays at the Monte Carlo rate.
The main benefit of our matrix jackknife approach is that it applies to very general nonlinear matrix algorithms, such as the randomized SVD and Nyström approximation. These nonlinear algorithms are used far more widely in practice because they produce errors comparable with the best low-rank approximation [11]. The bootstrap is not currently justified in this setting, whereas the jackknife is guaranteed to produce a reliable posterior variance estimate.
@@ 4.1 Nyström approximation
    Draw a test matrix $\boldsymbol{\Omega}\in\mathbb{K}^{d\times s}$ with iid isotropic columns.
@@
Algorithm 1 provides an implementation with $q=0$ with tricks to improve its numerical stability adapted from [12, 27]. For $q>0$, it may be necessary to introduce additional orthogonalization steps for reasons of numerical stability [23, Alg. 5.2].

### v4

@@ Abstract
Randomized matrix algorithms have become workhorse tools in scientific computing and machine learning. To use these algorithms safely in applications, they should be coupled with posterior error estimates to assess the quality of the output. To meet this need, this paper proposes two diagnostics: a leave-one-out error estimator for randomized low-rank approximations and a jackknife resampling method to estimate the variance of the output of a randomized matrix computation. Both of these diagnostics are rapid to compute for randomized low-rank approximation algorithms such as the randomized SVD and randomized Nyström approximation, and they provide useful information that can be used to assess the quality of the computed output and guide algorithmic parameter choices.
@@ 1 Introduction
In recent years, randomness has become an essential tool in the design of matrix algorithms [5, 11, 16, 17, 33], with randomized algorithms proving especially effective for matrix low-rank approximation. To use these algorithms safely in practice, they should be supported by posterior error estimates and other quality metrics that inform the user about the accuracy of the computational output.
@@
By using novel downdating formulas (see Eqs. 15 and 17 below), we can rapidly compute both of these estimators for widely used low-rank approximation methods such as the randomized SVD and randomized Nyström approximation. Our diagnostics are also sample-efficient in the sense that they require no additional matrix–vector products or other information beyond that used in the original algorithm. The speed and efficiency of these diagnostics make them compelling additions to workflows involving randomized matrix computation.
@@ 1.1 Leave-one-out error estimation
We begin by motivating our first diagnostic, a leave-one-out estimator for the error of a low-rank matrix approximation. For concreteness, we introduce this estimate in the context of Nyström approximation of positive-semidefinite (psd) matrices; see Section 2 for a more general setting
Suppose we want to approximate a psd matrix $\boldsymbol{A}\in\real^{d\times d}$, which we can only access by the matrix–vector product operation $\boldsymbol{\omega}\mapsto\boldsymbol{A}\boldsymbol{\omega}$. Using the matrix–vector product operation, we can compute the matrix–matrix product $\boldsymbol{A}\boldsymbol{\Phi}$ with a (random) matrix $\boldsymbol{\Phi}\in\real^{d\times s}$ with $s$ columns and form the Nyström approximation [9, 28, 32]
@@
where $\boldsymbol{\Omega}$ is populated with statistically independent standard Gaussian entries. The quality of the Nyström approximation improves with a higher approximation rank $s$ and number of subspace iteration steps $q$. Using the forthcoming Algorithm 1, we can compute $\boldsymbol{A}\langle\boldsymbol{\Phi}\rangle$ in the form of a compact eigenvalue decomposition:
@@
where $\boldsymbol{V}\in\real^{d\times s}$ has orthonormal columns and $\boldsymbol{\Lambda}\in\real_{+}^{s\times s}$ is diagonal. The cost of this procedure is $\mathcal{O}(qs)$ matrix–vector products with $\boldsymbol{A}$ and $\mathcal{O}(ds^{2})$ additional operations.
@@
What is the most efficient way to estimate the error $\left\|\boldsymbol{A}-\boldsymbol{A}\langle\boldsymbol{\Phi}\rangle\right\|_{\rm F}$?
Inspired by this question, this article proposes the leave-one-out error estimator, which provides an estimate of $\left\|\boldsymbol{A}-\boldsymbol{A}\langle\boldsymbol{\Phi}\rangle\right\|_{\rm F}$ using only the information already collected from $\boldsymbol{A}$ to form the Nyström approximation.
@@
$\boldsymbol{\Omega}\mapsto\boldsymbol{X}=\boldsymbol{X}(\boldsymbol{\Omega})\coloneqq\boldsymbol{A}\langle\boldsymbol{A}^{q}\boldsymbol{\Omega}\rangle.$
Let $\boldsymbol{\Omega}^{(j)}$ denote $\boldsymbol{\Omega}$ without its $j$th column. Introduce replicates $\boldsymbol{X}^{(1)},\ldots,\boldsymbol{X}^{(s)}$ by recomputing $\boldsymbol{X}$ with each column of $\boldsymbol{\Omega}$ left out in turn:
$\boldsymbol{X}^{(j)}=\boldsymbol{X}\big(\boldsymbol{\Omega}^{(j)}\big)\quad\text{for $j=1,2,\ldots,s$}.$
Letting $\boldsymbol{\omega}_{j}$ denote the $j$th column of $\boldsymbol{\Omega}$ and $\left\|\cdot\right\|$ denote the Euclidean norm, we define the leave-one-out error estimator
$\widehat{\operatorname{Err}}^{2}(\boldsymbol{X},\boldsymbol{A})\coloneqq\frac{1}{s}\sum_{i=1}^{s}\big\|\big(\boldsymbol{A}-\boldsymbol{X}^{(j)}\big)\boldsymbol{\omega}_{j}\big\|^{2}.$
As we show in Theorem 2.1, this error estimator is an unbiased estimator for the mean-square error of the rank-(s1)(s-1) Nyström approximation:
@@
Once the Nyström approximation has been computed, $\widehat{\operatorname{Err}}(\boldsymbol{X},\boldsymbol{A})$ is cheap to form, requiring at most $\mathcal{O}(ds^{2})$ additional operations (and only $\mathcal{O}(s^{3})$ operations if $q=0$). See Section 4.1 for details. The error estimate $\widehat{\operatorname{Err}}(\boldsymbol{X},\boldsymbol{A})$ requires no information about $\boldsymbol{A}$ beyond what is required to form the approximation $\boldsymbol{X}$.
A good point of comparison for the leave-one-out error estimator is provided by the Girard–Hutchinson norm estimate [16, §4.8]
(4) $\widehat{\operatorname{Err}}^{2}_{\rm GH}(\boldsymbol{X},\boldsymbol{A})=\frac{1}{t}\sum_{i=1}^{t}\left\|(\boldsymbol{A}-\boldsymbol{X})\boldsymbol{\nu}_{i}\right\|^{2}\approx\left\|\boldsymbol{A}-\boldsymbol{X}\right\|_{\rm F}^{2}.$
Here, $\boldsymbol{\nu}_{1},\ldots,\boldsymbol{\nu}_{t}$ are independent standard Gaussian test vectors. This estimator requires $t$ matrix–vector products with $\boldsymbol{A}$ (beyond those used to form the approximation $\boldsymbol{X}$), and the norm estimate becomes more accurate with larger $t$. The Girard–Hutchinson norm estimator is equivalent to the Girard–Hutchinson trace estimator [16, §4.2] applied to $(\boldsymbol{A}-\boldsymbol{X})^{*}(\boldsymbol{A}-\boldsymbol{X})$ and served as our inspiration for the leave-one-out error estimator. The leave-one-out estimator improves on the Girard–Hutchinson estimator as it requires no additional matrix–vector products with $\boldsymbol{A}$ to compute. In addition, the quality of the leave-one-out estimator automatically improves when the approximation rank $s$ increases, whereas the Girard–Hutchinson estimate only improves by using more matrix–vector products.
Figure 1: Leave-one-out error estimator. Approximation error $\operatorname{Err}(\boldsymbol{X},\boldsymbol{A})/\left\|\boldsymbol{A}\right\|_{\rm F}$ and normalized error estimate $\widehat{\operatorname{Err}}(\boldsymbol{X},\boldsymbol{A})/\left\|\boldsymbol{A}\right\|_{\rm F}$ (left) and relative error for the leave-one-out and Girard–Hutchinson error estimators (right) for Nyström approximation $\boldsymbol{X}$ to a psd kernel matrix $\boldsymbol{A}\in\real^{10^{4}\times 10^{4}}$. Error and error estimate were computed by an average of $1000$ trials. Error bars show one standard deviation.
Figure 1 demonstrates the leave-one-out error estimator. In this figure, we apply single-pass Nyström approximation ($q=0$) to approximate psd kernel matrix $\boldsymbol{A}\in\real^{10^{4}\times 10^{4}}$ formed from a random subsample of $10^{4}$ points from the QM9 dataset [22, 23] using approximation ranks $5\leq s\leq 150$. In the left panel, we plot the mean error
@@
and leave-one-out error estimate $\widehat{\operatorname{Err}}(\boldsymbol{X},\boldsymbol{A})$, estimated using $1000$ trials. We see that the estimate $\widehat{\operatorname{Err}}(\boldsymbol{X},\boldsymbol{A})$ closely tracks the true error. Moreover, the error estimate $\widehat{\operatorname{Err}}(\boldsymbol{X},\boldsymbol{A})$ is fast to compute, with the error estimate taking less than 1% of the total runtime to form. The right panel compares the mean relative error
$\text{mean relative error}=\operatorname{\mathbb{E}}\left[\frac{|\left\|\boldsymbol{A}-\boldsymbol{X}\right\|_{\rm F}-\mathrm{Est}|}{\left\|\boldsymbol{A}-\boldsymbol{X}\right\|_{\rm F}}\right],\quad\mathrm{Est}\in\left\{\widehat{\operatorname{Err}}(\boldsymbol{X},\boldsymbol{A}),\widehat{\operatorname{Err}}_{\rm GH}(\boldsymbol{X},\boldsymbol{A})\right\}$
for both the leave-one-out and Girard–Hutchinson error estimators. Following [29, §7.9], we use $t=10$ matrix–vector products for the Girard–Hutchinson estimator. For $s\geq 25$, the leave-one-out estimator is more accurate than Girard–Hutchinson and, for all values of $s$, the leave-one-out estimator is cheaper to form than Girard–Hutchinson, requiring just $\mathcal{O}(s^{3})$ operations and no additional matrix–vector products.
@@ 1.2.3 Jackknife variance estimation
Guarantees for this estimator are provided in Theorem 3.1. Algorithm 3 and LABEL:list:spectral_clustering provide pseudocode and a MATLAB implementation of Nyström-accelerated spectral clustering with the jackknife estimate $\operatorname{Jack}(\boldsymbol{X})$. The cost of forming the jackknife estimate $\operatorname{Jack}(\boldsymbol{X})$ is $\mathcal{O}(s^{2}d)$, much faster than the $\mathcal{O}(qsd^{2})$ total cost of Nyström-accelerated spectral clustering.
@@ 1.2.4 Numerical example
the mean and standard deviation for the standard deviation estimate $\operatorname{Jack}(\boldsymbol{X})$, the relative error $\operatorname{Err}(\boldsymbol{\widehat{A}},\boldsymbol{A})/\left\|\boldsymbol{A}\right\|_{\rm F}$ for the Nyström approximation $\boldsymbol{\widehat{A}}$, and the empirical success probability for spectral clustering.
Figure 2 shows the results. The estimate $\operatorname{Jack}(\boldsymbol{X})$ overestimates the standard deviation by a modest amount (a factor of five at most). The jackknife is not quantitatively sharp, but it is a reliable indicator of whether the variance is high or low.
The jackknife estimate $\operatorname{Jack}(\boldsymbol{X})$ can be used to determine whether the Nyström-based approximations to the eigenvectors are accurate enough for clustering task. At rank $s=50$, clustering is never performed correctly, and the jackknife estimate $\operatorname{Jack}(\boldsymbol{X})\approx 2$ is high. As the approximation rank $s$ is increased, clustering begins to succeed with higher and higher probability and the jackknife estimate $\operatorname{Jack}(\boldsymbol{X})$ decreases, indicating reduced variability.
The success of the jackknife estimate $\operatorname{Jack}(\boldsymbol{X})$ should be contrasted with the failure of the Nyström error $\operatorname{Err}(\boldsymbol{\widehat{A}},\boldsymbol{A})$ as a useful diagnostic for spectral clustering correctness. The Nyström error $\operatorname{Err}(\boldsymbol{\widehat{A}},\boldsymbol{A})$ decreases at a steady rate as $s$ is increased—unlike the jackknife estimate, the plot of $\operatorname{Err}(\boldsymbol{\widehat{A}},\boldsymbol{A})$ does not show an inflection point indicating the transition between clustering failure and success.
On the basis of these experiments, we propose two possible uses for the jackknife estimate in a spectral clustering workflow:
@@
These uses demonstrate the potential for jackknife variance estimation to be helpful in incorporating randomized low-rank approximation into general-purpose software.
@@ 2.2 Alternatives
A second alternative is the Girard–Hutchinson norm estimator Eq. 4, discussed in Section 1.1. The leave-one-out estimator improves on the Girard–Hutchinson estimator as the leave-one-out estimate does not require any additional matrix–vector products and automatically improves in quality as the number of vectors $s$ is increased.
@@ 3.5 Related work: bootstrap for randomized matrix algorithms
Bootstrap resampling [6, §5], a close relative of the jackknife, has seen several applications to matrix computations. An early use case was to provide confidence intervals for eigenvalues and eigenvectors of sample covariance matrices [26, §7.2].
A more recent line of work, led by Lopes and collaborators, applies bootstrap resampling to randomized matrix algorithms [13, 14, 15, 34]. The work closest to ours [13] uses the bootstrap to provide asymptotically sharp estimates of the error quantiles with regards to general error metrics for each singular value and singular vector computed by a Monte Carlo-type sketched SVD algorithm. In this special case, the bootstrap provides more fine-grained information than the matrix jackknife. Unfortunately, this sketched SVD is a poor computational method because its error decays at the Monte Carlo rate.
The main benefit of our matrix jackknife approach is that it is effective for very general matrix algorithms, such as the randomized SVD and Nyström approximation. These algorithms are used far more widely in practice because they achieve spectral accuracy, producing errors comparable with the best low-rank approximation [11]. As we demonstrate in Appendix SM2, a straightforward application of the bootstrap variance of Efron [6, §5.1] to the randomized SVD can produce standard deviation estimates which are incorrect by over four orders of magnitude.
@@ 4.1 Nyström approximation
    Draw a test matrix $\boldsymbol{\Omega}\in\mathbb{K}^{n\times s}$ with iid isotropic columns.
@@
Algorithm 1 provides an implementation with $q=0$ with tricks to improve its numerical stability adapted from [12, 28]. For $q>0$, it may be necessary to introduce additional orthogonalization steps for reasons of numerical stability [24, Alg. 5.2].